OpenStudy (anonymous):

Explain how to find the diagonal distance across a square with sides measuring 4 inches? Find the distance in simplified radical form.

OpenStudy (chaise):

Very very easily. If you're aware of Pythagoras theorem this is a breeze! a = sqrt(4^2+4^2) a = sqrt(16+16) a =5.65 inches :) The distance of the diagonal of a square with 4 inches for a side is 5.65 inches.